OpenStudy (anonymous):

a + 4c = 85 3a + 2c = 105 Solve the above equations simultaneously to find the admission cost of an adult and the admission cost of a child.

Directrix (directrix):

1) a + 4c = 85 2) 3a + 2c = 105 ---> Multiply every member of this equation by ( -2 )

Directrix (directrix):

1) a + 4c = 85 2) 3a + 2c = 105 -- by ( - 2) ==> -6a - 4c = -210 Then, add the "new" equation 2) to equation 1)

Directrix (directrix):

1) a + 4c = 85 2') -6a - 4c = -210 ------------------------ The "c" variable will add out. Solve for a. Then, back-substitute to get the value of c. @joshhhunt

OpenStudy (anonymous):

Thank you

Directrix (directrix):

Did you get the answers?

OpenStudy (anonymous):

@Directrix Sorry for posting so much I was sure how to do a reply

OpenStudy (anonymous):

@Directrix The answers I got are a = 65 c = 5

Directrix (directrix):

1) a + 4c = 85 2') -6a - 4c = -210 =============== -5a + 0c = -125 -5a = -125 a = 25

Directrix (directrix):

How did you get a = 65? I'll check my work. @joshhhunt

OpenStudy (anonymous):

I think I got mixed up

Directrix (directrix):

) a + 4c = 85 2') -6a - 4c = -210 =============== -5a + 0c = -125 Did you get this far? ---->

Directrix (directrix):

If a = 25 , then to get c, go to one of the two original equations. a + 4c = 85 and substitute 25 for a. So, 25 + 4c = 85 4c = 85 - 25 4c = 60 c = 15

OpenStudy (anonymous):

Yeah I had the wrong Idea, however I undertand now.

Directrix (directrix):

Okay.

OpenStudy (anonymous):

@Directrix Would you be able to help me out with one more question? For this task I have to create a question for the following: You have decided to make some cakes for a charity’s cake stall and need to determine the total costs and amounts of ingredients needed for these cakes. You wouldn't know what type of question i'd use would you?

Directrix (directrix):

Is this problem supposed to end up being solved by two equations in two unknowns? What are the instructions?

OpenStudy (anonymous):

Ok I will screenshot the whole thing however the only thing I really need help understanding is what type of question I should use.

Directrix (directrix):

You have decided to make some cakes for a charity's cake stall and need to determine the total costs and amounts of ingredients needed for these cakes. We are required to set out our answers using the following headings: Description (Clearly describe the scenario) Question (Clearly state what it is you are trying to solve) Method (Clearly explain in words and diagrams how you go about working out your solution) Solution (Show the mathematical steps involved in answering your question. Include written explanations where necessary). Is this it?

OpenStudy (anonymous):

That's it!

Directrix (directrix):

Nobody has worked it in such a way that it is clear what to do. Check this link: http://tinyurl.com/d5r8222

OpenStudy (anonymous):

Ok, thanks.

Directrix (directrix):

You could make cupcakes out of cake mix and chocolate frosting. Purchase the cake mix and frosting already prepared. Then, you just have to add water, pour the dough in the cupcake pan, cook, cool, and then frost and be done with it. All you have to buy is dough and frosting.

OpenStudy (anonymous):

So do you think the question will be a specific answer or it could just be something I can come up with like the profit at the end of the day based on numbers in the question?

Directrix (directrix):

I assume that this is a "story problem" and you get to make up what you like. Josh is preparing 100 cup cakes for the charity stall for a fundraiser for Josh's favorite charity. The cup cakes will be made of vanilla cake mix and the tops of 40 of the cup cakes will be left plain. The others will be frosted with chocolate. The cake mix for each cup cake costs 50 cents. The chocolate frosting for each cake costs 25 cents. How much money will Josh need to have to pay cash for the ingredients for the cup cakes? C = 40 * .50 + 60* (.50 + .25) C = ?